n) A constant current of 6.00 amp is passed through a solution containing NaCl(aq) .for 12.0 2 e minutes. During this time, the CI(aq) is oxidized via 2 CI(Clh+ and the resulting Cl(g) is collected. What is the volume of the Cl2(g), expressed in mL, if its pressure is 0.700 atm. and its temperature is 21.0°C? Assume 100% current efficiency, and assume that the gas is ideal. Data: R 8.3145 J/(K mol) -0.0821 L atm/(K mol), faraday 96,500 coulombs.
Answer #1

According to faradays law of electrolysis,

m = {Q/F} / {M/Z} m= mass deposited at electrode

Q = charge required

F= Faradays constant

M= molecular mass of ion

Z = valency factor

Q= i × t i = current

t = time of flow of current

m/M = { (i × t)/ F} / {1/Z}

n = {(i×t)/F} / (1/Z} n = no. of moles

{(6×12×60)/96500} / (1/1)

= 0.044 moles

Using ideal gas equation,

P × V = n × R × T P = 0.700 atm

n = 0.044 R = 0.0821 Latm K-1mol-1

T = 273 + 21 = 294 K

V = (n×R×T)/P

= (0.044 × 0.0821 × 294) / 0.700

= 1.52 ml

1.52 ml of Cl2 gas is liberated.